OpenStudy (anonymous):

g(x) = 1 - 3x - 9^2 find the discriminant and describe the nature of the solutions. does this follow into the quadratic formula since there is only one X value?

OpenStudy (anonymous):

I assume the last term is -9x^2. The discriminant = b^2 - 4ac, where a = coefficient of x^2, b = coefficient of x, and c is the constant Can you find the value of the discriminant?

OpenStudy (anonymous):

there is no X after the -9, so could -9 still be used as the A value in this equation?

OpenStudy (anonymous):

well, then retype the question's function g(x) again, please. As there seems you left out something somewhere. @cbhl

OpenStudy (anonymous):

g(x) = 1 - 3x - 9^2 The function is presented as so and the I am asked to find the discriminant. C would = 1 and B would = -3 but can we use -9 as A or is there more work that needs to be done here?

OpenStudy (anonymous):

A is the coefficient of x^2. But you dont have an x^2 term. Thats why I thought that it said -9x^2. Then there's an error in the problem. But I will assume that they forgot to include the x in -9x^2. Assuming that I am correct, the discriminant will be equal to what?

OpenStudy (anonymous):

the discriminant would be equal to 45, which means there would be two solutions in that case.

OpenStudy (anonymous):

Correct. And since 45 is not a perfect square, the two solutions will be real numbers and irrational numbers. (If the discriminant was 36, then it is a perfect square, and the roots will be real and rational numbers).

OpenStudy (anonymous):

A perfect square is something that can be multiplied by itself. 6 * 6 = 36, so 36 would be a perfect square. Sounds right, I guess the instructor forgot an X value in the question, thank you.

OpenStudy (anonymous):

Welcome. (Also, if the discriminant is negative, like -3, then the roots are imaginary (complex) numbers).

OpenStudy (anonymous):

Right, cheers.

OpenStudy (anonymous):

One final note. If the discriminant is 0, then the roots are equal and rational.

OpenStudy (anonymous):

There would be one solution, in that case.

OpenStudy (anonymous):

Yes, that is what's meant by "equal roots".